OpenStudy (anonymous):

Simplify: (x + 5)^2 A. x^2 + 25 B. x^2 + 5x + 25 C. x^2 + 10x + 25 D. x^2 + 10

OpenStudy (anonymous):

\[(x + 5)^2 = (x + 5)(x + 5)\]\[= x^2 + 5x + 5x + 25\]\[= x^2 + 10x + 25\]